OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 8/(x^2)+4

OpenStudy (anonymous):

f(x) = 8/x + 4 g(x) = x^2

OpenStudy (anonymous):

can you just explain that for me simply?

OpenStudy (anonymous):

well, you just need to pick any bit of the equation with x in it. so i could have picked 8/(x^2) and this is the inner function g(x). then make whatevers left the outer function f(x) = x + 4 in this case.

OpenStudy (anonymous):

ok so if I have y=2/(x^2) +3 following what you said: f(x)= 2/x +4 and g(x)=x^2 would be correct?